OpenStudy (anonymous):

Find the slope intercept equation of the line segment that passes through (1,3) and (3,-3)

OpenStudy (anonymous):

first find the gradient between the 2 p0ints... then just substitute to general linear line equation...

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

ok... first u know right the formula how to find the gradient between 2 points...??? that is : \[m = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] and then by substituting one of the coordinate of the p0int to the general linear line equation . that is \[y - y _{1}= m(x-x _{1})\] where your \[y _{1} and x _{1} \] is your coordinates that you want to substitute

OpenStudy (anonymous):

@oheneba do you get it...???

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Eda2012 so m= Y₂ -y₁/ X₂ -X₁ =-3

OpenStudy (anonymous):

@oheneba yes...ok u got it right...

OpenStudy (anonymous):

@Eda2012 where do I go from here?

OpenStudy (anonymous):

then u substitute one of the coordinate of the point to the general formula...

OpenStudy (anonymous):

where \[y-y _{1} = m(x-x _{1})\] your m = -3 your \[x _{1}\] and \[y_{1}\] is your coordinate of the point.

OpenStudy (anonymous):

so y-3=-3(x-1)?

OpenStudy (anonymous):

@Eda2012 so y-3=-3(x-1)?

OpenStudy (anonymous):

yeah....thats right... now u just simplify your answer and you've done..

OpenStudy (anonymous):

y=-3(x+2)

OpenStudy (anonymous):

@Eda2012 y=-3(x+2)

OpenStudy (anonymous):

y=-3x-6

OpenStudy (anonymous):

@Eda2012 y=-3(x+2 y=-3x-6 ?

OpenStudy (anonymous):

no... it suppose to be like this... \[y=-3(x+2)+3\] you cannot simply add or minus any number in the bracket....

OpenStudy (anonymous):

sorry my mistakes the equation suppose to be like this \[y=-3(x-1)+3\]

OpenStudy (anonymous):

but still you cannot simply add or minus any number in the bracket with the number that outside the bracket...

OpenStudy (anonymous):

Thanks. So this will be the final answer, right?

OpenStudy (anonymous):

no...you have to simply it...

OpenStudy (anonymous):

And how do I do that?

OpenStudy (anonymous):

-3x-3+3?

OpenStudy (anonymous):

yes...but you did a careless mistake in your answer just know... it suppose to be -3x+3+3

OpenStudy (anonymous):

-3x+3+3...lol.. Thanks =3x+6?

OpenStudy (anonymous):

yes!!! welcome~